# Tree Embedding in High Dimensions:

### v1

@@ 7 $r_{i}\leftarrow\beta/\Gamma\cdot w_{i}$
That is, for each bucket that intersects with $B(p,r)$, we choose the nearest point to $p$ as the representative point. It remains to verify that this definition of $\operatorname{rep}$ satisfies Definition˜3.6. Fix some $p\in P,i\in[m],r\geq 0$.
Verifying Item˜1.
Recall Item˜1 states that for each $x\neq y\in\operatorname{rep}_{i}(p,r)$, $\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)=\emptyset$. Consider some $x\neq y\in\operatorname{rep}_{i}(p,r)$. Suppose for the contrary that $\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)\neq\emptyset$, then there exists some $z\in\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)$, which implies that $\varphi(x)=\varphi(z)=\varphi(y)$. But by the definition of $\operatorname{rep}_{i}(p,r)$, at most one point of each bucket may belong to $\operatorname{rep}_{i}(p,r)$ which is a contradiction. This verifies Item˜1.
Verifying Item˜2.
Recall Item˜2 states that $\forall r^{\prime}\in(0,r)$, $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r)$. Observe that for each $r^{\prime}\in(0,r)$, we have $B(p,r^{\prime})\cap\operatorname{buks}_{i}(P)\subseteq B(p,r)\cap\operatorname{buks}_{i}(P)$. On the other hand, for sets $S,T\subseteq\mathbb{R}^{d}$ such that $S\subseteq T$, it is immediate that $\operatorname{NN}_{i}^{\operatorname{buks}}(p,S)\subseteq\operatorname{NN}_{i}^{\operatorname{buks}}(p,T)$. Then the item follows by substituting $S=B(p,r^{\prime})\cap\operatorname{buks}_{i}(P)$ and $T=B(p,r)\cap\operatorname{buks}_{i}(P)$ into this fact.
Verifying Item˜3.
Recall Item˜3 states that $\operatorname{rep}_{i}(p,r)\subseteq B(p,r)\cap\operatorname{buks}_{i}(P)$ and $\operatorname{buks}_{i}^{P}(\operatorname{rep}_{i}(p,r))=\operatorname{buks}_{i}^{P}(B(p,r))$. We first prove that $\operatorname{rep}_{i}(p,r)\subseteq B(p,r)\cap\operatorname{buks}_{i}(P)$. By definition, for each $x\in\operatorname{rep}_{i}(p,r)$, there exists $y\in B(p,r)\cap\operatorname{buks}_{i}(P)$ such that $x=\operatorname{NN}_{i}^{\operatorname{buks}}(p,y)$. Now, observe that $\operatorname{dist}(p,\operatorname{NN}_{i}^{\operatorname{buks}}(p,y))\leq\operatorname{dist}(p,y)$, which implies $\operatorname{dist}(p,x)\leq\operatorname{dist}(p,y)\leq r$, that is, $x\in B(p,r)$. On the other hand, using the definition of $\operatorname{NN}$ we have $x\in\operatorname{buks}_{i}(y)$, which means $\varphi_{i}(x)=\varphi_{i}(y)$. Besides, using $y\in\operatorname{buks}_{i}(P)$ there exists $z\in P$ such that $\varphi(x)=\varphi(y)=\varphi(z)$, which implies that $x\in\operatorname{buks}_{i}(P)$. Thus, $x\in B(p,r)\cap\operatorname{buks}_{i}(P)$.
@@
This finishes the proof of Lemma˜3.7. ∎
@@
In the remainder of the section, we assume $\{\operatorname{rep}_{i}(p,r)\}_{i,p,r}$ are representative sets. We then identify a key representative $\operatorname{rep}^{*}_{i}(p,r)\in\operatorname{rep}_{i}(p,r)$ (for $i,p,r$) in Definition˜3.8. Roughly speaking, this $\operatorname{rep}^{*}_{i}(p,r)\in\operatorname{rep}_{i}(p,r)$ is a representative point, coming from the bucket that contains a minimizer of the $\pi$-value within $\tilde{B}_{i}^{P}(p,r)$.
@@
By Item˜3 of Definition˜3.6, we already know that the label $\ell_{p}^{(i)}$, which takes the min $\pi$-value in $\tilde{B}^{P}_{i}(p,r_{i})$, can be equivalently evaluated within that of $\operatorname{rep}_{i}(p,r)$. This $\operatorname{rep}^{*}_{i}(p,r)$ in Definition˜3.6 precisely realizes the minimizer of $\pi$-value within $\operatorname{rep}_{i}(p,r)$:
@@
This fact can be formally implied by combining Definition˜3.8 with Item˜3 of Definition˜3.6. The technical reason to consider this $\operatorname{rep}^{*}$ which may not be in $P$, instead of the minimizer in $\tilde{B}^{P}(p,r)$ which is a subset of $P$, is that it enables us to utilize properties in Definition˜3.6.
We now turn to the proof of Lemma˜3.3.
Proof of Lemma˜3.3
@@
Specifically, we show that the event $\ell_{p}^{(i)}<\ell_{q}^{(i)}$, implies a geometric fact regarding the key representative $\operatorname{rep}^{*}$; specifically, $\operatorname{rep}^{*}_{i}(p,r_{i})\not\in B(q,r_{i})$. To see this, by (13) and Line 1 of Algorithm˜1, we have $\ell_{p}^{(i)}=\pi_{\min}(\operatorname{buks}_{i}^{P}(\operatorname{rep}^{*}_{i}(p,r_{i})))<\ell_{q}^{(i)}=\pi_{\min}(\operatorname{buks}^{P}_{i}(B(q,r_{i}))$, and this implies that $\operatorname{buks}_{i}^{P}(\operatorname{rep}_{i}^{*}(p,r_{i}))$ cannot be a subset of $\operatorname{buks}_{i}^{P}(B(q,r_{i}))$, which leads to $\operatorname{rep}_{i}^{*}(p,r_{i})\not\in B(q,r_{i})$. Therefore,
@@
We relate $\Pr[\operatorname{rep}^{*}_{i}(p,r_{i})\not\in B(q,r_{i})]$, which is what we need to prove as in (14), to the event $\mathcal{E}_{p}^{(i)}$ in the following Lemma˜3.10.
@@
where the first equality follows from the fact that $x=\operatorname{rep}_{i}^{*}(p,r_{i})\in\operatorname{rep}_{i}(p,r_{i})\subseteq B(p,r_{i})$ (by Item˜3 of Definition˜3.6), the second equality uses that $\mathcal{E}_{p}^{(i)}(\operatorname{rep}_{i}^{*}(p,r_{i}))$ always happens (which is immediate from Definition˜3.8 and Equation˜15), and the last equality holds because $\mathcal{E}_{p}^{i}(x)$ is independent to the randomness of $r_{i}$. ∎
@@
where the equality follows from the fact that $\operatorname{rep}_{i}^{*}(p,r_{i})\in\operatorname{rep}_{i}(p,r_{i})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})$ (which uses Item˜2 of Definition˜3.6) and that the events $x=\operatorname{rep}_{i}^{*}(p,r_{i})$ for every $x\in\operatorname{rep}_{i}(p,r_{i}^{\max})$ are mutually exclusive, and the last inequality follows from Lemma˜3.10.
@@
where the second-last inequality uses that $t_{j-1}+b_{j}=t_{j}$, which follows from Item˜1 of Definition˜3.6 (so sets in $\{\operatorname{buks}_{i}^{P}(p_{t}^{(i)})\}_{t}$ are disjoint). Then $\Pr[\mathcal{E}_{p}^{(i)}(p_{j}^{(i)})]\leq 1\leq 2(H_{t_{j}}-H_{t_{j-1}})$ trivially holds.
@@
We start with $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$. We know $\operatorname{rep}_{i}(p,r^{\prime})\subseteq B(p,r^{\prime})\cap\operatorname{buks}_{i}(P)$ by Item˜3 of Definition˜3.6, and $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})$ by Item˜2 of Definition˜3.6 and $r^{\prime}\leq r_{i}^{\max}$. These imply one direction: $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$. For the other direction, by Item˜3 of Definition˜3.6,
@@
so for every $x\in\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$ there exists some $y\in\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})$ such that $\operatorname{buks}_{i}^{P}(x)=\operatorname{buks}_{i}^{P}(y)$. Since we already established that $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$, so we have $y\in\operatorname{rep}_{i}(p,r_{i}^{\max}\cap B(p,r^{\prime}))$ as well. Therefore, it must be the case that $x=y$, since other, by Item˜1 of Definition˜3.6, we would have $\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)=\emptyset$, $\operatorname{buks}_{i}^{P}(x)\cap\operatorname{buks}_{i}^{P}(y)=\emptyset$, and this is a contradiction. This concludes the other direction, and finishes the proof of the claim. ∎
This claim implies $s^{\prime}=|\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})|=|\operatorname{rep}_{i}(p,r^{\prime})|$. By Item˜2 of Definition˜3.6, we have $\{p_{1}^{(i)},\ldots,p_{s^{\prime}}^{(i)}\}=\operatorname{rep}_{i}(p,r^{\prime})$ (recalling that the full set $\{p_{1}^{(i)},\ldots,p_{s}^{(i)}\}=\operatorname{rep}_{i}(p,r_{i}^{\max})$). By the premises and Item˜2 of Definition˜3.6, we have $\operatorname{rep}_{i}(p,r^{\prime})\subseteq B(p,r^{\prime})\subseteq B(q,\frac{1}{2}r_{i}^{\max})$, and this implies $p_{j}^{(i)}\in B(p,\frac{1}{2}r_{i}^{\max})$ for each $j\leq s^{\prime}$. Then using $r_{i}\geq\frac{1}{2}r_{i}^{\max}$, we have $p_{j}^{(i)}\in B(p,r_{i})$, that is,
@@
This finishes the proof of Lemma˜3.3. $\square$
We now turn to the proof of Lemma˜3.5, used in the similar-distance analysis (Item˜2) of Lemma˜3.3.
@@
Proof of Lemma˜3.5.
Compared with Lemma˜3.3, the main difference in Lemma˜3.5 is that it is without the premises $B(p,r^{\prime})\subseteq B(q,\frac{1}{2}r_{i}^{\max})$. Observe that it is only near the end of the proof of Lemma˜3.3 where we use this premises, which is precisely to show (19). Hence, (16), (17), (18) are still valid for Lemma˜3.5, and we can obtain the following.
@@
This finishes the proof of Lemma˜3.5. ∎
4 Dynamic and MPC Implementations of Algorithm˜1
In this section we present how to implement Algorithm˜1 efficiently for Euclidean $\mathbb{R}^{d}$, in dynamic and MPC settings. The key is to implement Line 1, particularly to bound the size of $\tilde{B}$’s. To this end, we make use of the following notion of consistent hashing.

### v2

@@ 7 $r_{i}\leftarrow\beta/\Gamma\cdot w_{i}$
That is, for each bucket that intersects with $B(p,r)$, we choose the nearest point to $p$ as the representative point. It remains to verify that this definition of $\operatorname{rep}$ satisfies Definition 3.6. Fix some $p\in P,i\in[m],r\geq 0$.
Verifying Item 1.
Recall Item 1 states that for each $x\neq y\in\operatorname{rep}_{i}(p,r)$, $\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)=\emptyset$. Consider some $x\neq y\in\operatorname{rep}_{i}(p,r)$. Suppose for the contrary that $\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)\neq\emptyset$, then there exists some $z\in\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)$, which implies that $\varphi(x)=\varphi(z)=\varphi(y)$. But by the definition of $\operatorname{rep}_{i}(p,r)$, at most one point of each bucket may belong to $\operatorname{rep}_{i}(p,r)$ which is a contradiction. This verifies Item 1.
Verifying Item 2.
Recall Item 2 states that $\forall r^{\prime}\in(0,r)$, $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r)$. Observe that for each $r^{\prime}\in(0,r)$, we have $B(p,r^{\prime})\cap\operatorname{buks}_{i}(P)\subseteq B(p,r)\cap\operatorname{buks}_{i}(P)$. On the other hand, for sets $S,T\subseteq\mathbb{R}^{d}$ such that $S\subseteq T$, it is immediate that $\operatorname{NN}_{i}^{\operatorname{buks}}(p,S)\subseteq\operatorname{NN}_{i}^{\operatorname{buks}}(p,T)$. Then the item follows by substituting $S=B(p,r^{\prime})\cap\operatorname{buks}_{i}(P)$ and $T=B(p,r)\cap\operatorname{buks}_{i}(P)$ into this fact.
Verifying Item 3.
Recall Item 3 states that $\operatorname{rep}_{i}(p,r)\subseteq B(p,r)\cap\operatorname{buks}_{i}(P)$ and $\operatorname{buks}_{i}^{P}(\operatorname{rep}_{i}(p,r))=\operatorname{buks}_{i}^{P}(B(p,r))$. We first prove that $\operatorname{rep}_{i}(p,r)\subseteq B(p,r)\cap\operatorname{buks}_{i}(P)$. By definition, for each $x\in\operatorname{rep}_{i}(p,r)$, there exists $y\in B(p,r)\cap\operatorname{buks}_{i}(P)$ such that $x=\operatorname{NN}_{i}^{\operatorname{buks}}(p,y)$. Now, observe that $\operatorname{dist}(p,\operatorname{NN}_{i}^{\operatorname{buks}}(p,y))\leq\operatorname{dist}(p,y)$, which implies $\operatorname{dist}(p,x)\leq\operatorname{dist}(p,y)\leq r$, that is, $x\in B(p,r)$. On the other hand, using the definition of $\operatorname{NN}$ we have $x\in\operatorname{buks}_{i}(y)$, which means $\varphi_{i}(x)=\varphi_{i}(y)$. Besides, using $y\in\operatorname{buks}_{i}(P)$ there exists $z\in P$ such that $\varphi(x)=\varphi(y)=\varphi(z)$, which implies that $x\in\operatorname{buks}_{i}(P)$. Thus, $x\in B(p,r)\cap\operatorname{buks}_{i}(P)$.
@@
This finishes the proof of Lemma 3.7. ∎
@@
In the remainder of the section, we assume $\{\operatorname{rep}_{i}(p,r)\}_{i,p,r}$ are representative sets. We then identify a key representative $\operatorname{rep}^{*}_{i}(p,r)\in\operatorname{rep}_{i}(p,r)$ (for $i,p,r$) in Definition 3.8. Roughly speaking, this $\operatorname{rep}^{*}_{i}(p,r)\in\operatorname{rep}_{i}(p,r)$ is a representative point, coming from the bucket that contains a minimizer of the $\pi$-value within $\tilde{B}_{i}^{P}(p,r)$.
@@
By Item 3 of Definition 3.6, we already know that the label $\ell_{p}^{(i)}$, which takes the min $\pi$-value in $\tilde{B}^{P}_{i}(p,r_{i})$, can be equivalently evaluated within that of $\operatorname{rep}_{i}(p,r)$. This $\operatorname{rep}^{*}_{i}(p,r)$ in Definition 3.6 precisely realizes the minimizer of $\pi$-value within $\operatorname{rep}_{i}(p,r)$:
@@
This fact can be formally implied by combining Definition 3.8 with Item 3 of Definition 3.6. The technical reason to consider this $\operatorname{rep}^{*}$ which may not be in $P$, instead of the minimizer in $\tilde{B}^{P}(p,r)$ which is a subset of $P$, is that it enables us to utilize properties in Definition 3.6.
We now turn to the proof of Lemma 3.3.
Proof of Lemma 3.3.
@@
Specifically, we show that the event $\ell_{p}^{(i)}<\ell_{q}^{(i)}$, implies a geometric fact regarding the key representative $\operatorname{rep}^{*}$; specifically, $\operatorname{rep}^{*}_{i}(p,r_{i})\not\in B(q,r_{i})$. To see this, by (13) and Line 1 of Algorithm 1, we have $\ell_{p}^{(i)}=\pi_{\min}(\operatorname{buks}_{i}^{P}(\operatorname{rep}^{*}_{i}(p,r_{i})))<\ell_{q}^{(i)}=\pi_{\min}(\operatorname{buks}^{P}_{i}(B(q,r_{i}))$, and this implies that $\operatorname{buks}_{i}^{P}(\operatorname{rep}_{i}^{*}(p,r_{i}))$ cannot be a subset of $\operatorname{buks}_{i}^{P}(B(q,r_{i}))$, which leads to $\operatorname{rep}_{i}^{*}(p,r_{i})\not\in B(q,r_{i})$. Therefore,
@@
We relate $\Pr[\operatorname{rep}^{*}_{i}(p,r_{i})\not\in B(q,r_{i})]$, which is what we need to prove as in (14), to the event $\mathcal{E}_{p}^{(i)}$ in the following Lemma 3.10.
@@
where the first equality follows from the fact that $x=\operatorname{rep}_{i}^{*}(p,r_{i})\in\operatorname{rep}_{i}(p,r_{i})\subseteq B(p,r_{i})$ (by Item 3 of Definition 3.6), the second equality uses that $\mathcal{E}_{p}^{(i)}(\operatorname{rep}_{i}^{*}(p,r_{i}))$ always happens (which is immediate from Definition 3.8 and Equation 15), and the last equality holds because $\mathcal{E}_{p}^{i}(x)$ is independent to the randomness of $r_{i}$. ∎
@@
where the equality follows from the fact that $\operatorname{rep}_{i}^{*}(p,r_{i})\in\operatorname{rep}_{i}(p,r_{i})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})$ (which uses Item 2 of Definition 3.6) and that the events $x=\operatorname{rep}_{i}^{*}(p,r_{i})$ for every $x\in\operatorname{rep}_{i}(p,r_{i}^{\max})$ are mutually exclusive, and the last inequality follows from Lemma 3.10.
@@
where the second-last inequality uses that $t_{j-1}+b_{j}=t_{j}$, which follows from Item 1 of Definition 3.6 (so sets in $\{\operatorname{buks}_{i}^{P}(p_{t}^{(i)})\}_{t}$ are disjoint). Then $\Pr[\mathcal{E}_{p}^{(i)}(p_{j}^{(i)})]\leq 1\leq 2(H_{t_{j}}-H_{t_{j-1}})$ trivially holds.
@@
We start with $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$. We know $\operatorname{rep}_{i}(p,r^{\prime})\subseteq B(p,r^{\prime})\cap\operatorname{buks}_{i}(P)$ by Item 3 of Definition 3.6, and $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})$ by Item 2 of Definition 3.6 and $r^{\prime}\leq r_{i}^{\max}$. These imply one direction: $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$. For the other direction, by Item 3 of Definition 3.6,
@@
so for every $x\in\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$ there exists some $y\in\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})$ such that $\operatorname{buks}_{i}^{P}(x)=\operatorname{buks}_{i}^{P}(y)$. Since we already established that $\operatorname{rep}_{i}(p,r^{\prime})\subseteq\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})$, so we have $y\in\operatorname{rep}_{i}(p,r_{i}^{\max}\cap B(p,r^{\prime}))$ as well. Therefore, it must be the case that $x=y$, since other, by Item 1 of Definition 3.6, we would have $\operatorname{buks}_{i}(x)\cap\operatorname{buks}_{i}(y)=\emptyset$, $\operatorname{buks}_{i}^{P}(x)\cap\operatorname{buks}_{i}^{P}(y)=\emptyset$, and this is a contradiction. This concludes the other direction, and finishes the proof of the claim. ∎
This claim implies $s^{\prime}=|\operatorname{rep}_{i}(p,r_{i}^{\max})\cap B(p,r^{\prime})|=|\operatorname{rep}_{i}(p,r^{\prime})|$. By Item 2 of Definition 3.6, we have $\{p_{1}^{(i)},\ldots,p_{s^{\prime}}^{(i)}\}=\operatorname{rep}_{i}(p,r^{\prime})$ (recalling that the full set $\{p_{1}^{(i)},\ldots,p_{s}^{(i)}\}=\operatorname{rep}_{i}(p,r_{i}^{\max})$). By the premises and Item 2 of Definition 3.6, we have $\operatorname{rep}_{i}(p,r^{\prime})\subseteq B(p,r^{\prime})\subseteq B(q,\frac{1}{2}r_{i}^{\max})$, and this implies $p_{j}^{(i)}\in B(p,\frac{1}{2}r_{i}^{\max})$ for each $j\leq s^{\prime}$. Then using $r_{i}\geq\frac{1}{2}r_{i}^{\max}$, we have $p_{j}^{(i)}\in B(p,r_{i})$, that is,
@@
This finishes the proof of Lemma 3.3. ∎
We now turn to the proof of Lemma 3.5, used in the similar-distance analysis (Item 2) of Lemma 3.3.
@@
Proof of Lemma 3.5.
Compared with Lemma 3.3, the main difference in Lemma 3.5 is that it is without the premises $B(p,r^{\prime})\subseteq B(q,\frac{1}{2}r_{i}^{\max})$. Observe that it is only near the end of the proof of Lemma 3.3 where we use this premises, which is precisely to show (19). Hence, (16), (17), (18) are still valid for Lemma 3.5, and we can obtain the following.
@@
This finishes the proof of Lemma 3.5. ∎
4 Dynamic and MPC Implementations of Algorithm 1
In this section we present how to implement Algorithm 1 efficiently for Euclidean $\mathbb{R}^{d}$, in dynamic and MPC settings. The key is to implement Line 1, particularly to bound the size of $\tilde{B}$’s. To this end, we make use of the following notion of consistent hashing.